OpenStudy (anonymous):

Rationalize the denominator of 5/9+5√5 ??

OpenStudy (anonymous):

do you mean 5/(9+5sqrt(5))

OpenStudy (anonymous):

\[\frac{5}{9+5\sqrt{5}}=\frac{5}{9+5\sqrt{5}}*\frac{9-5\sqrt{5}}{9-5\sqrt{5}}\] \[=\frac{45-25\sqrt{5}}{81-125}\]\[=\frac{-45+25\sqrt{5}}{44}\]

OpenStudy (anonymous):

you could type it into your calculator to get the answer :)